A hiker travels 4,5 km in a direction 12 degrees south of west and then 7,2 km in a direction 23 degrees west of south.





Determine the hiker’s overall displacement.





Draw each of the hiker’s individual displacement vectors on the same set of Cartesian axes, so that their initial points coincide at

1
Expert's answer
2022-04-13T17:21:53-0400

"r_W=4 .5\cos12°+7.2\sin 23°=7.2~km,"

"r_S=4.5\sin12°+7.2\cos23°=7.6~km,"

"r=\sqrt{r_W^2+r_S^2}=10.4~km,"

"\beta=\arctan\frac{r_S}{r_W}=46.5°" S of W.